Find an equation for the tangent plane to the surface 2xz^2-3xy-4x=7 at the point (1, -1, 2).



Select one:



A. 1




B. 0


1
Expert's answer
2022-04-11T10:34:54-0400

In case if a surface is defined implicitly by an equation of the form "F(x,y,z)=0" , then the tangent plane to the surface at a point "(a,b,c)"  is given by the equation:

"\frac{\partial F(a,b,c)}{\partial x}(x−a)+\frac{\partial F(a,b,c)}{\partial y}(y−b)""+\frac{\partial F(a,b,c)}{\partial z}(z−c)=0"

For the function "F(x,y,z)=2xz^2-3xy-4x-7" we have

"\frac{\partial F}{\partial x}=2z^2-3y-4" ; "\frac{\partial F}{\partial y}=-3x" ; "\frac{\partial F}{\partial z}=4xz", so the equation of the tangent plane at (1, -1, 2) is

"(2\cdot2^2-3\cdot(-1)-4)(x−1)+(-3\cdot1)(y−(-1))""+(4\cdot 1\cdot 2)(z−2)=0"

"7(x\u22121)-3(y+1)+8(z\u22122)=0"

"7x-3y+8z-26=0"

Answer: "7x-3y+8z-26=0" .